\begin{thebibliography}{9}
    		
    		\bibitem{AkiAusNag17}
    		E. Akin, J. Auslander and A. Nagar, Dynamics of induced systems,
    		\textit{Ergodic Theory Dynam.\ Systems},  37 (2017),  7,  {2034}--{2059}.
    		
    		\bibitem{Anosov67}
    		D. V. Anosov, Geodesic flows on closed Riemannian manifolds with negative curvature,
    		\textit{Proc. Steklov Inst. Mat.}, 90 (1967), {1}--{235}.
    		
    		\bibitem{WBauKSig75}
    		W. Bauer and K. Sigmund, Topological dynamics of transformations induced on the space of probability measures,
    		\textit{Monatsh.\ Math.}, 79 (1975), {81}--{92}.
    		
    		\bibitem{NBerUDarRVer22}
    		N. C. Bernardes Jr.,  U. B. Darji  and  R. M. Vermersch, Uniformly positive entropy of induced transformations,
    		\textit{Ergodic Theory Dynam.\ Systems},  42 (2022), {9}--{18}.
    		
    		\bibitem{NBerRVer16}
    		N. C. Bernardes Jr.\ and R. M. Vermersch, On the dynamics of induced maps on the space of probability
    		measures,
    		\textit{Trans.\ Amer.\ Math.\ Soc.}, 368 (2016), 11, {7703}--{7725}.
    		
    		\bibitem{PBil99}
    		P. Billingsley, Convergence of Probability Measures, Second Edition,
    		John Wiley \& Sons, Inc., New York, 1999.
    		
    		\bibitem{Burguet_Shi}
    		D.Burguet and R. Shi, Topological mean dimension of induced systems,
    		\textit{Trans.\ Amer.\ Math.\ Soc.}, 378 (2025), {3085}--{3103}.
    		
    		\bibitem{Downarowicz_Huczek}
    		T.Downarowicz and D. Huczek, Faithful zero-dimensional principal extensions,
    		\textit{Studia Mathematica}, 212 (1) (2012).
    		
    		\bibitem{RDud02}
    		R. M. Dudley, Real Analysis and Probability,
    		Cambridge Studies in Advanced Mathematics, 74,
    		Cambridge University Press, Cambridge, 2002.
    		
    		\bibitem{EGla83}
    		E. Glasner, Quasi-factors in ergodic theory,
    		\textit{Israel J. Math.}, 45 (1983), {198}--{208}.
    		
    		\bibitem{EGlabook03}
    		E. Glasner, Ergodic Theory via Joinings,
    		Mathematical Surveys and Monographs, vol 101, 2003.
    		
    		\bibitem{EGlaBWei95}
    		E. Glasner and B. Weiss, Quasi-factors of zero entropy systems,
    		\textit{J. Amer.\ Math.\ Soc.}, 8 (1995), 3, {665}--{686}.
    		
    		\bibitem{EGlaBWei03}
    		E. Glasner and B. Weiss, Quasifactors of ergodic systems with positive entropy,
    		\textit{Israel\ J.\ Math.}, 134 (2003), {363}--{380}.		
    		
    		\bibitem{Katok07}
    		A. Katok, Fifty years of entropy in dynamics:1958-2007,
    		\textit{Journal of Modern Dynamics}, 1 (2007), 4,  {545}--{596}.
    		
    		\bibitem{Katok80}
    		A. Katok, Lyapunov exponents, entropy and periodic orbits for diffeomorphisms,
    		\textit{Publ.\ Math.\ Inst.\ Hautes Études Sci}, 51 (1980), {137}--{173}.
    		
    		\bibitem{AKec95}
    		A. S. Kechris, Classical Descriptive Set Theory,
    		Springer-Verlag, New York, 1995.
    		
    		\bibitem{KerrLi09}
    		D. Kerr and H. Li, Combinatorial independence in measurable dynamics,
    		\textit{Journal of Functional Analysis}, 256 (2009), {1341}--{1386}.
    		
    		\bibitem{Kolmogorov}
    		A. N. Kolmogorov, New metric invariants of transitive dynamical systems and automorphisms of Lebesgue spaces,
    		\textit{Dokl. Akad. Nauk. SSSR}, 119 (1958), {861}--{864}.
    		
    		\bibitem{HanfengLi_KairanLiu}
    		H. Li and K. Liu, Local entropy theory, combinatorics and local theory of Banach spaces, 
    		\textit{Preprint}, arXiv:2507.03338.
    		
    		\bibitem{LiOprWu17}
    		J. Li, P. Oprocha and X. Wu, Furstenberg families, sensitivity and the space of probability
    		measures,
    		\textit{Nonlinearity}, 30 (2017), 3, {987}--{1005}.
    		
    		\bibitem{LiYanYe15}
    		J. Li, K. Yan and X. Ye, Recurrence properties and disjointness on the induced spaces,
    		\textit{Discrete Contin.\ Dyn.\ Syst.}, 35 (2015), 3, {1059}--{1073}.
    		
    		
    		\bibitem{Ornstein75}
    		D. S. Ornstein, Ergodic Theory, Randomness and Dynamical Systems,
    		Yale Mathematical Monographs, 5, Yale University Press, New Haven, Connecticut, 1975.
    		
    		\bibitem{Parry}
    		W. Parry, Topics in ergodic theory,
    		Cambridge tracts in mathematics, 75,
    		Cambridge University Press, Cambridge, 1981.
    		
    		\bibitem{Pesin77}
    		Ya. B. Pesin, Charactheristic Lyapunov exponents and smooth ergodic theory,
    		\textit{Russian Math. Surveys}, 32, 4 (1977), {55}--{114}.
    		
    		\bibitem{Pinsker}
    		M. S. Pinsker, Dynamical systems with completely positive or zero entropy,
    		\textit{Dokl. Akad. Nauk. SSSR}, 133 (1960), {937}--{938}.
    		
    		\bibitem{QiaoZhou17}
    		Y. Qiao and X. Zhou, Zero sequence entropy and entropy dimension,
    		\textit{Discrete Contin.\ Dyn.\ Syst.}, 37 (2017), 1, {435}--{448}.
    		
    		\bibitem{Rudolph}
    		D. Rudolph, Fundamentals of Measurable Dynamics,
    		Clarendon Press, Oxford, 1990.
    		
    		\bibitem{ShiZhang}
    		R. Shi and G. Zhang, Mean topological dimension of induced amenable group actions, \textit{Preprint}, arXiv:2507.03338.
    		
    		\bibitem{Shields}
    		P. C. Shields, The Ergodic Theory of Discrete Sample Paths,
    		Graduate Texts in Mathematics,13, American Mathematical Society, 1996.
    		
    		\bibitem{KSig78}
    		K. Sigmund, Affine transformations on the space of probability measures,
    		\textit{Ast\'erisque}, 51 (1978), {415}--{427}.
    		
    		\bibitem{Smale}
    		S. Smale, Differentiable dynamical systems,
    		\textit{Bull.\ Amer.\ Math.\ Soc.}, 73 (1967), {747}--{817}.
    		
    		\bibitem{Vermersch}
    		R. M. Vermersch, Measure-theoretic uniformly positive entropy on the space of probability measures,
    		\textit{Qual. Theory Dyn. Syst.}, 23 (2024), 1, 29.
    		
    		\bibitem{Yomdin}
    		Y. Yomdin, Volume growth and entropy,
    		\textit{Israel\ J.\ Math.}, 57 (1987), {285}--{300}.
    		
    		\bibitem{Young}
    		L-S. Young, Dimension, entropy and Lyapunov exponents,
    		\textit{Ergodic Theory Dynam.\ Systems}, 2 (1982), 1, {109}--{124}.
    		
    		
    	\end{thebibliography}

\documentclass[12pt,oneside,reqno]{amsart}

\usepackage[utf8]{inputenc}
\usepackage[a4paper,nomarginpar]{geometry}

\usepackage[english]{babel}
\usepackage{enumitem}
\usepackage{amsmath,amsfonts,amssymb,amsthm}
\usepackage{amscd}
\usepackage{tikz}

\allowdisplaybreaks

\makeatletter
\def\subsection{\@startsection{subsection}{2}%
	\z@{.5\linespacing\@plus.7\linespacing}{.3\linespacing}%
	{\normalfont\bfseries}}
\makeatother

\newtheorem{theorem}{Theorem}%[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

\theoremstyle{definition}

\begin{document}
	
	% \title[short text for running head]{full title}
	
	\title[On the entropy of processes generated by quasifactors]
	{On the entropy of processes generated by quasifactors}
	
	% Only \author and \address are required; other information is optional.
	% Remove any unused author tags.
	
	% author one information
	% \author[short version for running head]{name for top of paper}

    \author{R\^omulo M. Vermersch}
	\address{Departamento de Matem\'atica, Centro de Ci\^encias F\'isicas e
		Matem\'aticas, Universidade Federal de Santa Catarina,
		Florian\'opolis, SC, 88040-900, Brazil.}
	\curraddr{}
	\email{romulo.vermersch@ufsc.br}
	\thanks{}

	% author two information
	
%	\author{}
%	\address{}
%	\curraddr{}
%	\email{}
%	\thanks{}

	% \subjclass is required.
	
	\subjclass[2010]{Primary 28D05; Secondary 28A33, 28D20, 60B10.}
	\keywords{entropy, homeomorphisms,
		probability measures, quasifactors.}
	
	\date{}
	
	\dedicatory{}
	\begin{abstract}
		
		Given a measurable dynamical system $(X,\mathcal{X},\mu,T)$ with $h_{\mu}(T)<\infty$, where $X$ is a compact metric space, $\mathcal{X}$ is the Borel $\sigma$-algebra on $X$, $\mu$ is a $T$-invariant Borel probability measure and $T$ is a homeomorphism acting on $X$ we show that, if $h_{\mu}(T)>0$, then $h_{\widetilde{\mu}}(\widetilde{T})>0$ for every quasifactor $\widetilde{\mu}$ of $\mu$ having full-support.
	\end{abstract}
    
	\maketitle

	%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

	\section{Introduction}\label{Intro}

    Entropy is a central notion in ergodic theory, providing a fundamental measure of the unpredictability and complexity of a dynamical system. Actually, since Kolmogorov's foundational work \cite{Kolmogorov}, entropy has become a major theme within important directions of research such as: isomorphism theory \cite{Ornstein75}, Lyapunov exponents \cite{Katok80,Pesin77,Young}, volume growth rates \cite{Yomdin}, uniformly \cite{Anosov67,Smale} and non-uniformly \cite{Pesin77} hyperbolic dynamical systems. A detailed account of the deep connections between entropy and these topics can be found in the expository paper by Katok \cite{Katok07}, where many more references can also be found. 
    
    \smallskip
    
    By a {\em measurable dynamical system} (MDS) we mean a quadriple $\mathfrak{X}=(X,\mathcal{X},\mu,T)$, where $X$ is a compact metric space, $\mathcal{X}$ is the Borel $\sigma$-algebra on $X$, $\mu$ is a Borel probability measure on $\mathcal{X}$ and $T:X\to X$ is a homeomorphism that preserves $\mu$.
    
    By a {\em topological dynamical system} (TDS) we mean a pair $(X,T)$
    consisting of a compact metric space $X$ and a homeomorphism $T : X \to X$. 
    Such a TDS induces, in a natural way, the TDS $(\mathcal{M}(X),\widetilde{T}).$
    Here, $\mathcal{M}(X)$ denotes the space of all Borel probability
    measures on $X$ endowed with the {\em Prokhorov metric}
    $$
    d_P(\mu,\nu):= \inf\{\delta > 0 : \mu(A) \leqslant \nu(A^\delta) + \delta
    \text{ for all } A \in \mathcal{X}\},
    $$
    and
    $\widetilde{T} : \mathcal{M}(X) \to \mathcal{M}(X)$ is the homeomorphism given by
    $$
    (\widetilde{T}(\mu))(A):= \mu(T^{-1}(A)) \ \ \ (\mu \in \mathcal{M}(X), A \in \mathcal{X}).
    $$
    
    It is well known that  $\mathcal{M}(X)$ is a compact metric space and that $d_P(\mu,\nu)$ induces the so-called
    {\em weak*-topology} on $\mathcal{M}(X)$, that is, the topology whose basic open neighborhoods
    of $\mu \in \mathcal{M}(X)$ are the sets of the form
    $$
    \mathbb{V}(\mu;f_1,\ldots,f_k;\varepsilon):= \Big\{\nu \in \mathcal{M}(X) :
    \Big| \int_X f_i \,d\nu - \int_X f_i \,d\mu \Big| < \varepsilon
    \text{ for } i = 1,\ldots,k\Big\},
    $$
    where $k \geqslant 1$, $f_1,\ldots,f_k : X \to \mathbb{R}$ are continuous functions and
    $\varepsilon > 0$.
    
    We refer the reader to the books
    \cite{PBil99,RDud02,AKec95} for a study of the space $\mathcal{M}(X)$.
    
    \smallskip

    The research on the connections between the dynamics of the TDS $(X,T)$ and
    the dynamics of the induced TDS $(\mathcal{M}(X),\widetilde{T})$ was initiated by Bauer and Sigmund \cite{WBauKSig75},
    and was later developed by several authors; see \cite{NBerUDarRVer22,NBerRVer16,Burguet_Shi,EGlaBWei95,EGlaBWei03,LiOprWu17,LiYanYe15,QiaoZhou17,ShiZhang,KSig78}, for instance.
    The TDS $(\mathcal{M}(X),\widetilde{T})$ serves as an abstract model for systems in statistical mechanics, where the dynamics can be described in deterministic terms (in the sense that the time-evolution of the system is given by some physical law), but the states of the system are probability distributions on the phase space. 
    In \cite{EGla83} Glasner introduced the notion of a {\em quasifactor} of a MDS as an ergodic-theoretic analogue of an induced TDS. Let us see its definition:
    
    A {\it quasifactor} of $\mathfrak{X}=(X,\mathcal{X},\mu,T)$ is a MDS $\widetilde{\mathfrak{X}}=(\mathcal{M}(X),\widetilde{\mathcal{X}},\widetilde{\mu},\widetilde{T})$ such that  $\widetilde{\mu}$ satisfies the so-called {\it barycenter equation}:
    \begin{equation}\label{barycenter} 
    	\mu=\displaystyle\int_{\mathcal{M}(X)}\theta d\widetilde{\mu}(\theta)
    \end{equation}
    Here, $\widetilde{\mathcal{X}}$ denotes the Borel $\sigma$-algebra on $\mathcal{M}(X)$.
    Equivalently, we say that {\em $\mu$ is the barycenter of $\widetilde{\mu}$}.
    
    The barycenter equation means that, by choosing any compact topology on $X$ compatible with its Borel structure one has
    \begin{equation*}\displaystyle\int_{X}f(x)d \mu(x)=\int_{\mathcal{M}(X)}\int_{X}f(x)d \theta(x)d \widetilde{\mu}(\theta)
    \end{equation*}
    for all $f:X\rightarrow\mathbb{R}$ continuous function.
    
    Glasner also showed that this definition is independent of the choice of the compact topology compatible with the Borel structure (\cite{EGla83}). 
    For convenience, sometimes we will say that $\widetilde{\mu}$ is a quasifactor of $\mu$ and we shall denote by $Q(\mu)$ the set of all quasifactors of $\mu$.
    
    Moreover we remark that, for each fixed $A\in\mathcal{X}$ the map $\Psi_{A}:\nu\in\mathcal{M}(X)\mapsto \nu(A)\in [0,1]$ is Borel and $\displaystyle \mu(A)=\int_{\mathcal{M}(X)}\nu(A)d \widetilde{\mu}(\nu)$; for a proof of this well-known fact see Lemma 4.1 from \cite{HanfengLi_KairanLiu}.
    
    \smallskip
    
    In this work we are concerned with the relationship between the entropy of the MDS $\mathfrak{X}=(X,\mathcal{X},T,\mu)$ and of the MDS $\widetilde{\mathfrak{X}}=(\mathcal{M}(X),\widetilde{\mathcal{X}},\widetilde{T},\widetilde{\mu})$, where $\widetilde{\mu}\in Q(\mu)$.

    The research on the relationship between the entropy of a MDS and of a quasifactor of it can be traced back to a deep result due to Glasner and Weiss \cite{EGlaBWei95} which asserts that
    if $\mathfrak{X}=(X,\mathcal{X},T,\mu)$ has zero entropy, then so does $\widetilde{\mathfrak{X}}=(\mathcal{M}(X),\widetilde{\mathcal{X}},\widetilde{T},\widetilde{\mu})$ for {\em every} $\widetilde{\mu}\in Q(\mu)$. By the variational principle it implies that, if $(X,T)$ has topological zero entropy, then so does $(\mathcal{M}(X),\widetilde{T})$. We mention that Qiao and Zhou \cite{QiaoZhou17} obtained such a result for the
    notion of sequence entropy.
    
    In another work, Glasner and Weiss \cite{EGlaBWei03} proved that any ergodic system of positive entropy admits {\em every} ergodic system of positive entropy as a quasifactor, which shows, in particular, that the set of quasifactors of an ergodic system of positive entropy is very large.
    
    We also mention that in \cite{Vermersch} the author initiated the investigation on the relationship between the entropy of the MDS $\mathfrak{X}=(X,\mathcal{X},T,\mu)$ and $\widetilde{\mathfrak{X}}=(\mathcal{M}(X),\widetilde{\mathcal{X}},\widetilde{T},\widetilde{\mu})$ in the context of local entropy theory \cite{KerrLi09}. Very recently, Li and Liu, among other findings, expanded it and extended it to amenable group actions \cite{HanfengLi_KairanLiu}.
    \smallskip
    
    Let $A\in\mathcal{X}$ be such that $0<\mu(A)<1$, $\mu(\partial A)=0$ and $\mu\Big(\displaystyle\overline{\bigcup_{r\in\mathbb{Z}}\partial T^{r}A}\Big)=0$. Notice that, if $X$ is a zero-dimensional space and $A\in\mathcal{X}$ is any clopen set, then $\displaystyle\overline{\bigcup_{r\in\mathbb{Z}}\partial T^{r}A}=\varnothing$ and so, we have $\mu\Big(\displaystyle\overline{\bigcup_{r\in\mathbb{Z}}\partial T^{r}A}\Big)=0$. Now, let $0<\eta<1$ and put $\widetilde{A}=\{\nu\in\mathcal{M}(X):\nu(A)>\eta\}$. Since $A\in\mathcal{X}$, $\Psi_{A}(\nu)=\nu(A)$ is a Borel map and $\widetilde{A}=\Psi^{-1}(\eta,1]$, we see that $\widetilde{A}\in\widetilde{\mathcal{X}}$. Write $\mathcal{P}=\{A,A^{c}\}$ and $\widetilde{\mathcal{P}}=\{\widetilde{A},\widetilde{A}^{c}\}$. So, $\mathcal{P}$ is a non-trivial two-set partition of $X$ into Borel sets and, as we shall see, if $\widetilde{\mu}\in Q(\mu)$ has full-support (i.e. if it is positive on the non-empty open sets of $\mathcal{M}(supp\ \mu)$), then $\widetilde{\mathcal{P}}$ is a non-trivial two-set partition of $\mathcal{M}(X)$ into Borel sets (Proposition~\ref{twosetpartition}). Observe that, if $supp\ \mu\neq X$, then it is not possible to find a quasifactor of $\mu$ having full-support on all $\mathcal{M}(X)$. So, when saying that a quasifactor has full-support, it is understood that it has full-support on $\mathcal{M}(supp\ \mu)$. It turns out that our main result (Theorem~\ref{Mainresult}) is based on an analysis of the relationship between the entropy of the stationary stochastic processes generated by the pairs $(\mathfrak{X},\mathcal{P})$ and $(\widetilde{\mathfrak{X}},\widetilde{\mathcal{P}})$, where $\widetilde{\mu}\in Q(\mu)$ has full-support. In fact, we shall show that, if $\widetilde{\mu}\in Q(\mu)$ has full-support and $(\widetilde{\mathfrak{X}},\widetilde{\mathcal{P}})$ has zero entropy, then $(\mathfrak{X},\mathcal{P})$ has zero entropy (Theorem~\ref{zeroentropy}). In addition, if $\mu$ and $\widetilde{\mu}$ are {\em ergodic}, then we show that this fact occurs {\em continuously} (Theorem~\ref{ergodic_case}). We begin our analysis by the ergodic case, taking advantage of the characterization of the entropy of an ergodic finite-valued stochastic process in terms of the {\em covering-exponent} property to show the aforementioned continuity property. In the case where $\mu$ and $\widetilde{\mu}$ are not necessarily ergodic, we prove that by showing that the present of the process $(\mathfrak{X},\mathcal{P})$ can be arbitrarily well predicted from its past, given that the present of the process $(\widetilde{\mathfrak{X}},\widetilde{\mathcal{P}})$ is sufficiently predictable from its past (Theorem~\ref{approximatedbypast}). As a consequence, by using the Downarowicz-Huczek theorem about zero-dimensional extensions \cite{Downarowicz_Huczek} we obtain our main result: if $0<h_{\mu}(T)<\infty$, then $h_{\widetilde{\mu}}(\widetilde{T})>0$ for every $\widetilde{\mu}\in Q(\mu)$ having full-support (Theorem~\ref{Mainresult}). We remark that we cannot omit the full-support hypothesis for $\widetilde{\mu}\in Q(\mu)$ even in the ergodic case. Actually, we can have $h_{\mu}(T)>0$ and, if we consider $\widetilde{\mu}:=\delta_{\mu}$, then $\widetilde{\mu}\in Q(\mu)$, $\widetilde{\mu}$ is ergodic and $h_{\widetilde{\mu}}(\widetilde{T})=0$.

    \section[Preliminaries]{Preliminaries}
    
    Let us recall some definitions and notation from entropy theory.
    In what follows, all logarithms are in base $e$. 
    
    Let $\mathfrak{X}=(X,\mathcal{X},\mu,T)$ be a MDS.
    Given a finite partition $\mathcal{P}=\{P_{0},P_{1},\ldots,P_{k-1}\}$ of $X$, we consider the so-called {\em name map} $\Phi_{\mathcal{P}}:X\to\{0,1,\ldots,k-1\}^{\mathbb{Z}}$ defined by:
    $$(\Phi_{\mathcal{P}}(x))_{n}=j\ \text{if and only if}\  T^{n}x\in P_{j}\ \ (0\leqslant j\leqslant k-1, n\in\mathbb{Z}).
    $$
    The sequence $(\Phi_{\mathcal{P}}(\cdot))_{n\in\mathbb{Z}}$ is a {\em stationary stochastic process}. We say that $(\Phi_{\mathcal{P}}(\cdot))_{n\in\mathbb{Z}}$ is {\em the process generated by $\mathfrak{X}$ and the partition $\mathcal{P}$}.
    
    If $f$ is a random variable in $X$ taking values in $\{0,\ldots,k-1\}$ and we consider, for each $0\leqslant j \leqslant k-1$, the set $P_{j}:=f^{-1}(\{j\})$, then we see that $\mathcal{P}:=\{P_{0},\ldots,P_{k-1}\}$ is a partition of $X$ into Borel sets. Hence, since we can think of a finite partition as a finite-valued random variable that assigns to each point the set containing it, we obtain a correspondence between finite partitions and finite-valued random variables.
    
    The {\em entropy} of a random variable $f$ associated with the finite partition $\mathcal{P}$ is defined by
    
    $$
    H(f):= -\sum_{P\in\mathcal{P}}\mu(P)\log \mu(P),
    $$
    We also write $H(\mathcal{P})=H(f)$.

    Given a stochastic process $(f_{i})_{i\in\mathbb{Z}}$ on $X$ taking values in the finite set $\{0,\ldots,k-1\}$, for each $n\geqslant 1$ we define the {\em joint} of $f_{0},\ldots,f_{n-1}$ by: $$\displaystyle\bigvee_{i=0}^{n-1}f_{i}:=\{P_{0} \cap \cdots \cap P_{n-1}:
    P_{0} \in \mathcal{P}_{0},\ldots,P_{n-1} \in \mathcal{P}_{n-1}\}=\bigvee_{i=0}^{n-1}\mathcal{P}_{i},$$
    where $\mathcal{P}_{i}$ is the partition of $X$ corresponding to $f_{i}$ ($0\leqslant i\leqslant n-1$).

    The {\em entropy of the stochastic process $(f_{i})_{i\in\mathbb{Z}}$} is defined by the following expression:
    $$H((f_{i})_{-\infty}^{+\infty}):= \displaystyle\lim_{n \to \infty}(1/n) H\big(\bigvee_{i=0}^{n-1}\mathcal{P}_{i}\big).
    $$
    
    The {\em entropy of $T$ with respect to $\mathcal{P}$} is defined by
    $$
    h_{\mu}(T,\mathcal{P}):= \displaystyle\lim_{n \to \infty}(1/n) H\big(\bigvee_{i=0}^{n-1}T^{-i}\mathcal{P}\big).
    $$
    Clearly, we have $h_{\mu}(T,\mathcal{P})=H((\Phi_{\mathcal{P}})_{-\infty}^{+\infty})$.
    
    Given $n\in\mathbb{N}$, a finite partition $\mathcal{P}$ of $X$ into Borel sets and $0<\gamma<1$ we denote by $N(n,\mathcal{P},T,\gamma)$ the minimum cardinality of a subcollection $\displaystyle\mathcal{G}\subseteq\bigvee_{i=0}^{n-1}T^{-i}\mathcal{P}$ needed to cover a Borel set $D\subseteq X$ with $\mu(D)\geqslant\gamma$. If $(X,\mathcal{X},\mu,T)$ is \emph{ergodic}, then $h_{\mu}(T,\mathcal{P})$ has the {\em covering-exponent} property; that is,
    
    $$h_{\mu}(T,\mathcal{P})=\displaystyle\lim_{n\to\infty}(1/n)\log N(n,\mathcal{P},T,\gamma)\ \text{for every}\ 0<\gamma<1 $$\ (see, for example, Theorem 5.1 on page 72 from \cite{Rudolph} or Theorem I.7.4 on page 68 from \cite{Shields}).

    Finally, the {\em entropy} of $T$ is given by
    $$
    h_{\mu}(T):= \sup_{\mathcal{P}} h_{\mu}(T,\mathcal{P}),
    $$
    where the supremum is taken over all finite partitions $\mathcal{P}$ of $X$ into Borel sets.
    
    In other words, the entropy of $T$ is the supremum over all the entropies of processes of form $(\Phi_{\mathcal{P}})$ with $\mathcal{P}$ being a finite partition of $X$ into Borel sets.
    
    Let $\Pi\subseteq\mathcal{X}$ be the smallest $\sigma$-algebra containing the collection of all sets $A\in\mathcal{X}$ with $h_{\mu}(T,\{A,A^{c}\})=0$. Pinsker \cite{Pinsker} defined $\Pi$ and showed that:
    
    \begin{itemize}
    	\item[(i)]$T^{-1}\Pi =\Pi$;
    	\item[(ii)]If $\mathcal{F}$ is a $T$-invariant $\sigma$-algebra such that $\mathcal{F}\subseteq\Pi$, then $h_{\mu}(T,\{A,A^{c}\})=0$ for every $A\in\mathcal{F}$.
    \end{itemize}
    Thus, $\Pi$ is the largest $T$-invariant $\sigma$-algebra ``with'' zero entropy. We call $\Pi$ the {\em Pinsker $\sigma$-algebra} of the dynamical system $\mathfrak{X}$. Furthermore, we call the restriction of the dynamical system $\mathfrak{X}$ to $\Pi$ the {\em Pinsker factor} of $\mathfrak{X}$. The Pinsker factor is the {\em deterministic part} of $\mathfrak{X}$. The books by Glasner \cite{EGlabook03} and by Parry \cite{Parry} are standard references for the study of the Pinsker factor.
    Note that $h_{\mu}(T)=0$ if, and only if, $\Pi=\mathcal{X}$. Equivalently, we have $h_{\mu}(T)>0$ if, and only if, there exists some $A\in\mathcal{X}$ with $h_{\mu}(T,\{A,A^{c}\})>0$.
    
    We denote by $\displaystyle\bigvee_{i=0}^{\infty}T^{-i}\mathcal{P}$ the smallest complete $\sigma$-algebra containing all atoms of $\displaystyle\bigvee_{i=0}^{n-1}T^{-i}\mathcal{P}$ for every $n\geqslant 1$. Since $\displaystyle\bigvee_{i=0}^{n-1}T^{-i}\mathcal{P}\subset\bigvee_{i=0}^{n}T^{-i}\mathcal{P}$ for every $n\geqslant 1$ we write $\displaystyle\bigvee_{i=0}^{n-1}T^{-i}\mathcal{P}\uparrow\bigvee_{i=0}^{\infty}T^{-i}\mathcal{P}$
    
    \section{Our results}
    
    We begin with a simple result concerning two-set partitions:
    
    \begin{proposition}{\label{twosetpartition}} Let $(X,\mathcal{X},\mu,T)$ be a MDS, $0<\mu(A)<1$, $\mu(\partial A)=0$, $0<\eta<1$ and let $\widetilde{\mu}\in Q(\mu)$ having full-support. If $\widetilde{A}:=\{\nu\in\mathcal{M}(X):\nu(A)>\eta\}$, then $0<\widetilde{\mu}(\widetilde{A})<1$. 
    	\begin{proof} Suppose $\widetilde{\mu}(\widetilde{A})=0$. So, $\widetilde{\mu}(\{\nu:\nu(A)\leqslant \eta\})=1$. On the other hand, since $\mu(\partial A)=0$, it follows from the barycenter equation that $\nu(\partial A)=0$ for $\widetilde{\mu}$-a.e. $\nu\in\mathcal{M}(X)$. Therefore, $\nu(\overline{A})=\nu(A)$ for $\widetilde{\mu}$-a.e. $\nu\in\mathcal{M}(X)$. So, we get $\widetilde{\mu}(\{\nu:\nu(\overline{A})\leqslant \eta\})=1$. Since $\widetilde{\mu}(\mathcal{M}(supp\ \mu))=1$, we obtain $\widetilde{\mu}(\{\nu:\nu(\overline{A})\leqslant\eta\}\cap\mathcal{M}(supp\ \mu))=1$. Hence, since $\widetilde{\mu}$ has full-support, it follows that $\{\nu:\nu(\overline{A})\leqslant \eta\}\cap\mathcal{M}(supp\ \mu)$ is dense in $\mathcal{M}(supp\ \mu)$. But, since $\{\nu:\nu(\overline{A})\leqslant \eta\}\cap\mathcal{M}(supp\ \mu)$ is closed in $\mathcal{M}(supp\ \mu)$, we obtain $\mathcal{M}(supp\ \mu)=\{\nu:\nu(\overline{A})\leqslant \eta\}\cap\mathcal{M}(supp\ \mu)$, which is impossible. The same argument can be used to show that we cannot have $\widetilde{\mu}(\widetilde{A})=1$ either. Hence, we obtain $0<\widetilde{\mu}(\widetilde{A})<1$. This completes the proof.
    		
    	\end{proof}
    \end{proposition}
    
    Let $A$ and $\widetilde{A}$ be as above and put $P_{0}:=A$, $P_{1}:=A^{c}$, $\widetilde{P}_{0}:=\widetilde{A}$, $\widetilde{P}_{1}:=\widetilde{A}^{c}$. So, $\mathcal{P}:=\{P_{0},P_{1}\}$ is a non-trivial partition of $X$ into Borel sets and, by Proposition~\ref{twosetpartition} we see that $\widetilde{\mathcal{P}}:=\{\widetilde{P}_{0},\widetilde{P}_{1}\}$ is a non-trivial partition of $\mathcal{M}(X)$ into Borel sets. To avoid unnecessary repetitions, for the rest of the paper we shall consider $A\in\mathcal{X}$ with $0<\mu(A)<1$, $\mu(\partial A)=0$, $0<\eta<1$, $\widetilde{A}=\{\nu\in\mathcal{M}(X):\nu(A)>\eta\}$ and write $\mathcal{P}=\{P_{0},P_{1}\}$ and $\widetilde{\mathcal{P}}=\{\widetilde{P}_{0},\widetilde{P}_{1}\}$, where $P_{0}=A$, $P_{1}=A^{c}$, $\widetilde{P}_{0}=\widetilde{A}$, $\widetilde{P}_{1}=\widetilde{A}^{c}$. We also put $N:=\displaystyle\overline{\bigcup_{r\in\mathbb{Z}}\partial T^{r}A}$ and suppose $\mu(N)=0$.
    
    \begin{lemma}\label{largemeasurelemma}Let $(X,\mathcal{X},\mu,T)$ be a MDS and let $\widetilde{\mu}\in Q(\mu)$ having full-support. 
    	
    	Given any $0<\gamma<1$ there exists $\gamma^{\prime}>0$  such that, for any $m,n\geqslant 1$ and any functions $\sigma_{1},\ldots,\sigma_{m}\in\{0,1\}^{\{0,\ldots,n-1\}}$, if $\displaystyle\widetilde{\mu}\Big(\bigcup_{l=1}^{m}\bigcap_{i=0}^{n-1}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{l}(i)}\Big)\geqslant \gamma^{\prime}$, then $\displaystyle\mu\Big(\bigcup_{l=1}^{m}\bigcap_{i=0}^{n-1}T^{-i}P_{\sigma_{l}(i)}\Big)\geqslant \gamma$. %equivalently, if $\displaystyle\widetilde{\mu}\Big(\bigcap_{l=1}^{m}\bigcup_{i=0}^{n-1}\widetilde{T}^{-i}\widetilde{P}_{\tau_{l}(i)}\Big)\geqslant 1-\gamma^{\prime}$, then $\displaystyle\mu\Big(\bigcap_{l=1}^{m}\bigcup_{i=0}^{n-1}T^{-i}P_{\tau_{l}(i)}\Big)\geqslant 1-\gamma$, where $\tau_{l}(i):=\{0,1\}\setminus\{\sigma_{l}(i)\}$.
    	\begin{proof}To obtain a contradiction, suppose that the result does not hold. In this case, there exists some $0<\gamma_{0}<1$ such that, for every $k\geqslant 1$ there are $m_{k},n_{k}\geqslant 1$ and functions $\sigma_{1}^{k},\ldots,\sigma_{m_{k}}^{k}\in\{0,1\}^{\{0,\ldots,n_{k}-1\}}$ such that $\displaystyle\widetilde{\mu}\Big(\bigcup_{l=1}^{m_{k}}\bigcap_{i=0}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{l}^{k}(i)}\Big)>1- 2^{-k}$ but $\displaystyle\mu\Big(\bigcup_{l=1}^{m_{k}}\bigcap_{i=0}^{n_{k}-1}T^{-i}P_{\sigma_{l}^{k}(i)}\Big)< \gamma_{0}$ for every $k\geqslant 1$. Put $\tau_{l}^{k}(i):=\{0,1\}\setminus\{\sigma_{l}^{k}(i)\}$. Since $\displaystyle\sum_{k=1}^{\infty}\displaystyle\widetilde{\mu}\Big(\bigcap_{l=1}^{m_{k}}\bigcup_{i=0}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{P}_{\tau_{l}^{k}(i)}\Big)< \infty$, it follows from the Borel-Cantelli lemma that , for $\widetilde{\mu}$-a.e. $\nu\in\mathcal{M}(X)$ there exists some $k_{0}=k_{o}(\nu)\geqslant 1$ such that $\nu\notin \displaystyle \bigcap_{l=1}^{m_{k}}\bigcup_{i=0}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{P}_{\tau_{l}^{k}(i)}$ for all $k\geqslant k_{0}$. In other words, if  
    		$$\widetilde{\mathcal{G}}:=\Big\{\nu\in\mathcal{M}(X):\exists k_{0}\geqslant 1; \nu\in\displaystyle\bigcup_{l=1}^{m_{k}}\bigcap_{i=1}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{l}^{k}(i)}\ \text{for all}\ k\geqslant k_{0}\Big\},$$
    		then $\widetilde{\mu}(\widetilde{\mathcal{G}})=1$. %So, by the barycenter equation we see that there exists $\widetilde{\mathcal{K}}\subset\mathcal{M}(X)$ with $\widetilde{\mu}(\widetilde{\mathcal{K}})=1$ such that $\nu(N)=0$ for every $\nu\in\widetilde{\mathcal{K}}$. 
    		Since $\widetilde{\mu}(\widetilde{\mathcal{G}}\cap\mathcal{M}(supp\ \mu))=1$ and $\widetilde{\mu}$ has full-support, it follows that $\widetilde{\mathcal{G}}\cap\mathcal{M}(supp\ \mu)$ is dense in $\mathcal{M}(supp\ \mu)$. Now, since $\displaystyle\mu\Big(\bigcap_{l=1}^{m_{k}}\bigcup_{i=0}^{n_{k}-1}T^{-i}P_{\tau_{l}^{k}(i)}\Big)>1- \gamma_{0}$ for all $k\geqslant 1$ and $\mu(N)=0$, we may pick some $x\in supp\ \mu$, $x\notin N$ such that $x\in\displaystyle\bigcap_{l=1}^{m_{k}}\bigcup_{i=1}^{n_{k}-1}T^{-i}P_{\tau_{l}^{k}(i)}$ for infinitely many $k's$. Consequently, $\delta_{x}\in\mathcal{M}(supp\ \mu)$ is such that $\delta_{x}\in\displaystyle\bigcap_{l=1}^{m_{k}}\bigcup_{i=1}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{P}_{\tau_{l}^{k}(i)}$ for infinitely many $k's$. So, since $\widetilde{\mathcal{G}}\cap\mathcal{M}(supp\ \mu)$ is dense in $\mathcal{M}(supp\ \mu)$, for every $\varepsilon>0$ there are $k_{0}\geqslant 1$ and $\nu^{\prime}\in\mathcal{M}(supp\ \mu)$ satisfying:
    		\begin{align}&\nu^{\prime}\in\displaystyle\bigcup_{l=1}^{m_{k}}\bigcap_{i=1}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{l}^{k}(i)} \ \text{for all}\ k\geqslant k_{0}\label{nuprimeinLemma}\ \text{and}\\%\\
    			%&\nu^{\prime}(N)=0\label{nuprimenullLemma}\ \text{and}\\
    			&\nu^{\prime}(B)\leqslant \delta_{x}(B^{\varepsilon})+\varepsilon \  \text{for all Borel sets} \ B\subseteq X\label{nuprimedeltaxLemma}.
    		\end{align}
    		
    		Now, fix some $k^{\prime}\geqslant k_{0}$ such that $\delta_{x}\in\displaystyle\bigcap_{l=1}^{m_{k^{\prime}}}\bigcup_{i=1}^{n_{k^{\prime}}-1}\widetilde{T}^{-i}\widetilde{P}_{\tau_{l}^{k^{\prime}}(i)}$. So, for every $1\leqslant l\leqslant m_{k^{\prime}}$ there exists some $0\leqslant i^{\prime}\leqslant n_{k^{\prime}}-1$ such that $\widetilde{T}^{i^{\prime}}\delta_{x}\in\widetilde{P}_{\tau_{l}^{k^{\prime}}(i^{\prime})}$. On the other hand, by \eqref{nuprimeinLemma} we see that there exists some $1\leqslant l^{\prime}\leqslant m_{k^{\prime}}$  such that $\widetilde{T}^{i}\nu^{\prime}\in\widetilde{P}_{\sigma_{l^{\prime}}^{k^{\prime}}(i)}$ for every $0\leqslant i\leqslant n_{k^{\prime}}-1$. In particular, for $l=l^{\prime}$ there exists some $0\leqslant i^{\prime}\leqslant n_{k^{\prime}}-1$ such that $\delta_{T^{i^{\prime}}x}(A)>\eta$ and $\nu^{\prime}(T^{-i^{\prime}}A^{c})\geqslant 1-\eta$ if $\tau_{l^{\prime}}^{k^{\prime}}(i^{\prime})=0$, or $\delta_{T^{i^{\prime}}x}(A^{c})\geqslant 1-\eta$ and $\nu^{\prime}(T^{-i^{\prime}}A)>\eta$ if $\tau_{l^{\prime}}^{k^{\prime}}(i^{\prime})=1$. Now, without loss of generality we may assume that $\tau_{l^{\prime}}^{k^{\prime}}(i^{\prime})=0$. In this case, we have $T^{i^{\prime}}x\in A$ and $\nu^{\prime}(T^{-i^{\prime}}A^{c})\geqslant 1-\eta$. Moreover, by \eqref{nuprimedeltaxLemma} with $B=T^{-i^{\prime}}A^{c}$ we get $\delta_{x}((T^{-i^{\prime}}A^{c})^{\varepsilon})\geqslant \nu^{\prime}(T^{-i^{\prime}}A^{c})-\varepsilon\geqslant 1-\eta-\varepsilon>0$, whenever $\varepsilon>0$ is small enough depending on $0<\eta<1$. Therefore, we see that $x\in T^{-i^{\prime}}A$ and $x\in [(T^{-i^{\prime}}A)^{c}]^{\varepsilon}$ for every $\varepsilon>0$ small enough depending on $0<\eta<1$. Finally, since $i^{\prime}=i^{\prime}(\varepsilon)$ we have to consider two cases:
    		\begin{itemize}
    			\item[(i)]The set $\{i^{\prime}(\varepsilon):\varepsilon>0\}$ is bounded as $\varepsilon\rightarrow 0$.
    			In this case, there are $i^{\prime}\geqslant 1$ and a sequence $\varepsilon_{n}\rightarrow 0$ such that $i^{\prime}=i^{\prime}(\varepsilon_{n})$ for all $n\geqslant 1$. So, we get $x\in T^{-i^{\prime}}A$ and $x\in [(T^{-i^{\prime}}A)^{c}]^{\varepsilon_{n}}$ for all $n\geqslant 1$. Therefore, by letting $n\rightarrow\infty$ we conclude that $x\in T^{-i^{\prime}}A\ \cap\overline{(T^{-i^{\prime}}A)^{c}}$, which contradicts the choice $x\notin N$.
    			\item[(ii)]The set $\{i^{\prime}(\varepsilon):\varepsilon>0\}$ is unbounded as $\varepsilon\rightarrow 0$. In this case, by letting $\varepsilon\rightarrow 0$ we obtain $\displaystyle x\in\overline{\bigcup_{r\in\mathbb{Z}}T^{r}A}\ \cap \overline{\bigcup_{r\in\mathbb{Z}}\partial T^{r}A^{c}}$, which contradicts the choice $x\notin N$ again.
    		\end{itemize} This proves the lemma.
    	\end{proof}
    \end{lemma}
    
    \smallskip
    
    By analogous reasoning, one can prove the following result:
    
    \begin{lemma}Let $(X,\mathcal{X},\mu,T)$ be a MDS and let $\widetilde{\mu}\in Q(\mu)$ having full-support. 
    	
    	Given any $0<\gamma<1$ there exists $\gamma^{\prime}>0$  such that, for any $m,n\geqslant 1$ and any functions $\sigma_{1},\ldots,\sigma_{m}\in\{0,1\}^{\{0,\ldots,n-1\}}$, if $\displaystyle\widetilde{\mu}\Big(\bigcup_{l=1}^{m}\bigcap_{i=0}^{n-1}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{l}(i)}\Big)\leqslant \gamma^{\prime}$, then $\displaystyle\mu\Big(\bigcup_{l=1}^{m}\bigcap_{i=0}^{n-1}T^{-i}P_{\sigma_{l}(i)}\Big)\leqslant \gamma$.
    \end{lemma}
    \smallskip
    
    \begin{theorem}\label{coveringtheorem}Let $(X,\mathcal{X},\mu,T)$ be a MDS and let $\widetilde{\mu}\in Q(\mu)$ having full-support. Given $\alpha>0$ there exists $\beta>0$ with the following property: 
    	
    	Given $0<\gamma<1$ there exist $\gamma^{\prime}>0$ and $n_{0}\geqslant 1$ such that, if $n\geqslant n_{0}$ and $N(n,\widetilde{\mathcal{P}},\widetilde{T},\gamma^{\prime})<e^{n\beta}$, then $N(n,\mathcal{P},T,\gamma)<e^{n\alpha}$.
    	\begin{proof}To obtain a contradiction, let us assume that the conclusion does not hold. So, there exists some $\alpha>0$ such that, for $\beta_{k}:=1/k$ there exists $0<\gamma_{\beta_{k}}=\gamma_{k}<1$ such that, for every $\gamma^{\prime}>0$ there exists $n_{k,\gamma^{\prime}}\geqslant k$ such that $N(n_{k,\gamma^{\prime}},\widetilde{\mathcal{P}},\widetilde{T},\gamma^{\prime})<e^{n_{k,\gamma^{\prime}}.k^{-1}}$ and $N(n_{k,\gamma^{\prime}},\mathcal{P},T,\gamma_{k})\geqslant e^{n_{k,\gamma^{\prime}}.\alpha}$\ ($k\geqslant 1$). Let $\gamma^{\prime}_{k}>0$ be associated to $\gamma_{k}$ according to Lemma~\ref{largemeasurelemma} and put $n_{k,\gamma^{\prime}_{k}}=n_{k}$. For every $k\geqslant 1$ we have:

    		\begin{equation}\label{ineq.1} N(n_{k},\widetilde{\mathcal{P}},\widetilde{T},\gamma^{\prime}_{k})<e^{n_{k}.k^{-1}}
    		\end{equation}
    		and
    		\begin{equation}\label{ineq.2}N(n_{k},\mathcal{P},T,\gamma_{k})\geqslant e^{n_{k}.\alpha}.
    		\end{equation}
    		Fix $k\geqslant 1$ large enough so that $1/k\leqslant \alpha/2$. Note that (\ref{ineq.1}) means that there exists $\widetilde{D}\subseteq\mathcal{M}(X)$ with $\widetilde{\mu}(\widetilde{D})\geqslant \gamma^{\prime}_{k}$ that admits a collection $\displaystyle\widetilde{\mathcal{G}}\subseteq\bigvee_{i=0}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{\mathcal{P}}$ with $|\widetilde{\mathcal{G}}|<e^{n_{k}.k^{-1}}$ as a cover $(*)$. Furthermore, note that \eqref{ineq.2} means that for every $D\subseteq X$ with $\mu(D)\geqslant \gamma_{k}$, if $\displaystyle\mathcal{G}\subseteq\bigvee_{i=0}^{n_{k}-1}T^{-i}\mathcal{P}$ is a collection that covers $D$, then $|\mathcal{G}|\geqslant e^{n_{k}.\alpha}$ $(**)$. Let $\Sigma$ be the collection of functions $\sigma:\{0,\ldots,n_{k}-1\}\rightarrow \{0,1\}$ such that, given any $\widetilde{G}\in\widetilde{\mathcal{G}}$ there exists a necessarily unique $\sigma\in\Sigma$ with $\widetilde{G}=\displaystyle\bigcap_{i=0}^{n_{k}-1}\widetilde{T}^{-i}\widetilde{P}_{\sigma(i)}$. We now define $\mathcal{G}\subseteq\displaystyle\bigvee_{i=0}^{n_{k}-1}T^{-i}\mathcal{P}$ as follows: $G\in\mathcal{G}$ if, and only if, there exists a (necessarily unique) $\sigma\in\Sigma$ such that $G=\displaystyle\bigcap_{i=0}^{n_{k}-1}T^{-i}P_{\sigma(i)}$. Put $D:=\displaystyle\bigcup_{G\in\mathcal{G}}G$. Clearly, $\mathcal{G}$ covers $D$ and $|\mathcal{G}|=|\widetilde{\mathcal{G}}|$. Moreover, since $\displaystyle\widetilde{\mu}\big(\bigcup_{\widetilde{G}\in\widetilde{\mathcal{G}}}\widetilde{G}\big)\geqslant \gamma^{\prime}_{k}$, it follows from Lemma~\ref{largemeasurelemma} that $\mu(D)\geqslant \gamma_{k}$. Therefore, from $(*)$ and $(**)$ we get:
    		$$e^{n_{k}.\alpha}\leqslant |\mathcal{G}|= |\widetilde{\mathcal{G}}|\leqslant e^{n_{k}.k^{-1}},
    		$$
    		which contradicts the choice $1/k\leqslant \alpha/2$. This concludes the proof of the theorem.
    	\end{proof}
    \end{theorem}
    
    \begin{theorem}\label{ergodic_case}Let $(X,\mathcal{X},\mu,T)$ be a MDS and let $\widetilde{\mu}\in Q(\mu)$ having full-support. If $\mu$ and $\widetilde{\mu}$ are ergodic, then the following continuity property holds:
    	
    	Given $\alpha>0$ there exists $\beta>0$ such that, if $h_{\widetilde{\mu}}(\widetilde{T},\widetilde{\mathcal{P}})<\beta$, then $h_{\mu}(T,\mathcal{P})<\alpha$.
    	\begin{proof} Let $\alpha>0$ be given and take $\beta>0$ as in Theorem~\ref{coveringtheorem}. Suppose $h_{\widetilde{\mu}}(\widetilde{T},\widetilde{\mathcal{P}})<\beta$. By Theorem~\ref{coveringtheorem}, given $0<\gamma<1$ there exist $\gamma^{\prime}>0$ and $n_{0}\geqslant 1$ such that, if $n\geqslant n_{0}$ and $N(n,\widetilde{\mathcal{P}},\widetilde{T},\gamma^{\prime})<e^{n.\beta}$, then $N(n,\mathcal{P},T,\gamma)<e^{n.\alpha/2}$. .Moreover, there exists $n^{\prime}_{0}\geqslant 1$ such that $N(n,\widetilde{\mathcal{P}},\widetilde{T},\gamma^{\prime})<e^{n.\beta}$ whenever $n\geqslant n^{\prime}_{0}$. So, if $n\geqslant\max\{n_{0},n^{\prime}_{0}\}$, then we have $n\geqslant n_{0}$ and $N(n,\widetilde{\mathcal{P}},\widetilde{T},\gamma^{\prime})<e^{n.\beta}$. Thus, we see that $N(n,\mathcal{P},T,\gamma)<e^{n.\alpha/2}$, whenever $n\geqslant\max\{n_{0},n^{\prime}_{0}\}$. Hence, we obtain $h_{\mu}(T,\mathcal{P})=\displaystyle\lim_{n\to\infty}(1/n)N(n,\mathcal{P},T,\gamma)\leqslant\alpha/2<\alpha$, as desired.
    	\end{proof}
    	
    \end{theorem}
    
    \medskip
    
    Now we turn to the case where both $\mu$ and $\widetilde{\mu}$ are not necessarily ergodic. For this end we need to recall that, given two finite partitions $\mathcal{P}$ and $\mathcal{Q}$ and given any $\varepsilon>0$ we write $\mathcal{P}\subseteq_{\varepsilon}^{\mu}\mathcal{Q}$ to mean that for every $P\in\mathcal{P}$ there exists some union $\bigcup Q$ of atoms of $\mathcal{Q}$ with $P\subseteq \bigcup Q$ and $\mu(\bigcup Q\setminus P)<\varepsilon$. Finally, we write $\mathcal{P}\subseteq_{0}^{\mu}\mathcal{Q}$ if, for every $P\in\mathcal{P}$ there exists some union $\bigcup Q$ of atoms of $\mathcal{Q}$ with $P\subseteq \bigcup Q$ and $\mu(\bigcup Q\setminus P)=0$. Of course, we have $\mathcal{P}\subseteq_{0}^{\mu}\mathcal{Q}$ if and only if $\mathcal{P}\subseteq_{\varepsilon}^{\mu}\mathcal{Q}$ for every $\varepsilon>0$.
    
    \smallskip
    The ideas behind the proof of our next result parallel with those of Lemma~\ref{largemeasurelemma}; in particular, the use of the classical Borel-Cantelli lemma is crucial again, though the required analysis is more involved.
    
    \begin{theorem}{\label{approximatedbypast}} Let $(X,\mathcal{X},\mu,T)$ be a MDS and let $\widetilde{\mu}\in Q(\mu)$ having full-support. 
    	
    	Given $0<\alpha<1$ there are $\beta>0$ and $n_{0}\geqslant 1$ such that, if $n\geqslant n_{0}$ and $\displaystyle\widetilde{\mathcal{P}}\subseteq_{\beta}^{\widetilde{\mu}}\bigvee_{i=1}^{n}\widetilde{T}^{-i}\widetilde{\mathcal{P}}$, then $\displaystyle\mathcal{P}\subseteq_{\alpha}^{\mu}\bigvee_{i=1}^{n}T^{-i}\mathcal{P}$.
    	
    	\begin{proof} Assume that the result does not hold. In this case, there exist $0<\alpha<1$ and an increasing sequence $n_{k}\rightarrow\infty$ with the following property:
    		\begin{equation}\label{widetildeapproximatedincluded}\displaystyle\widetilde{\mathcal{P}}\subseteq_{2^{-k}}^{\widetilde{\mu}}\bigvee_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{\mathcal{P}}
    		\end{equation}
    		
    		but
    		\begin{equation}\label{notapproximatedincluded}\displaystyle\mathcal{P}\subsetneq_{\alpha}^{\mu}\bigvee_{i=1}^{n_{k}}T^{-i}\mathcal{P}\  \textrm{for every}\ k\geqslant 1.
    		\end{equation}
    		Observe that (\ref{widetildeapproximatedincluded}) means that:
    		
    		For each $j\in\{0,1\}$ there are $q_{j,k}\geqslant 1$ and functions $\sigma_{j,k}^{1},\ldots,\sigma_{j,k}^{q_{j,k}}:\{1,\ldots,n_{k}\}\rightarrow\{0,1\}$ such that $\displaystyle\widetilde{P_{j}}\subseteq\bigcup_{l=1}^{q_{j,k}}\bigcap_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{j,k}^{l}(i)}$ and $\displaystyle\widetilde{\mu}\Big(\bigcup_{l=1}^{q_{j,k}}\bigcap_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{j,k}^{l}(i)}\setminus \widetilde{P}_{j}\Big)<2^{-k}$  ($k\geqslant 1$).
    		
    		Furthermore, observe that (\ref{notapproximatedincluded}) means that: 
    		
    		There exists $j^{\prime}\in\{0,1\}$ such that $\displaystyle\mu\Big(\bigcup_{l=1}^{q}\bigcap_{i=1}^{n_{k}}T^{-i}P_{\sigma^{l}(i)}\setminus P_{j^{\prime}}\Big)\geqslant\alpha$, whenever $q\geqslant 1$ and $\sigma^{1},\ldots,\sigma^{q}:\{1,\ldots,n_{k}\}\rightarrow\{0,1\}$ satisfy $\displaystyle P_{j^{\prime}}\subseteq\bigcup_{l=1}^{q}\bigcap_{i=1}^{n_{k}}T^{-i}P_{\sigma^{l}(i)}$.
    		
    		Fix $j=j^{\prime}$. Since $\displaystyle\sum_{k=1}^{\infty}\widetilde{\mu}\Big(\bigcup_{l=1}^{q_{j^{\prime},k}}\bigcap_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{j^{\prime},k}^{l}(i)}\setminus \widetilde{P}_{j^{\prime}}\Big)<+\infty$, it follows from the Borel-Cantelli lemma that, for $\widetilde{\mu}$-a.e. $\nu\in\mathcal{M}(X)$ there exists some $k_{0}=k_{0}(\nu)\geqslant 1$ such that $\displaystyle\nu\notin\bigcup_{l=1}^{q_{j^{\prime},k}}\bigcap_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{j^{\prime},k}^{l}(i)}\setminus\widetilde{P}_{j^{\prime}}$ for all $k\geqslant k_{0}$. That is, if $\tau_{j^{\prime},k}^{l}(i):=\{0,1\}\setminus \{\sigma_{j^{\prime},k}^{l}(i)\}$ and 
    		$$\widetilde{\mathcal{G}}:=\Big\{\nu\in\mathcal{M}(X):\exists k_{0}\geqslant 1; \nu\in\displaystyle\bigcap_{l=1}^{q_{j^{\prime},k}}\bigcup_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{P}_{\tau_{j^{\prime},k}^{l}(i)}\cup P_{j^{\prime}}\ \text{for all}\ k\geqslant k_{0}\Big\},$$
    		then $\widetilde{\mu}(\widetilde{\mathcal{G}})=1$.
    		Let us consider the following sets: $\widetilde{Q}_{0}:=\{\nu:\nu(\overline{A})\geqslant\eta\}$ and 
    		$\widetilde{Q}_{1}:=\{\nu:\nu(\overline{A^{c}})\geqslant1-\eta\}$. Thus, $\widetilde{Q}_{0}$ and $\widetilde{Q}_{1}$ are closed sets with $\widetilde{P}_{0}\subseteq\widetilde{Q}_{0}$ and $\widetilde{P}_{1}\subseteq\widetilde{Q}_{1}$. Let us consider the following set:
    		$$\widetilde{\mathcal{H}}:=\Big\{\nu\in\mathcal{M}(X):\exists k_{0}\geqslant 1; \nu\in\displaystyle\bigcap_{l=1}^{q_{j^{\prime},k}}\bigcup_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{Q}_{\tau_{j^{\prime},k}^{l}(i)}\cup Q_{j^{\prime}}\ \text{for all}\ k\geqslant k_{0}\Big\}.$$
    		Clearly, $\widetilde{\mathcal{G}}\subseteq\widetilde{\mathcal{H}}$ and so, $\widetilde{\mu}(\widetilde{\mathcal{H}})=1$.
    		Since $\mu(N)=0$, it follows from the barycenter equation that there exists $\widetilde{\mathcal{K}}\subset\mathcal{M}(X)$ with $\widetilde{\mu}(\widetilde{\mathcal{K}})=1$ such that $\nu(N)=0$ for every $\nu\in\widetilde{\mathcal{K}}$. Hence, $\widetilde{\mu}(\widetilde{\mathcal{H}}\cap\widetilde{\mathcal{K}}\cap\mathcal{M}(supp\ \mu))=1$.  Since $\widetilde{\mu}$ has full-support, it follows that $\widetilde{\mathcal{H}}\cap\widetilde{\mathcal{K}}\cap\mathcal{M}(supp\ \mu)$ is dense in $\mathcal{M}(supp\ \mu)$. Now, let us consider the set 
    		$$\widetilde{\Lambda}:=\Big\{\nu\in\mathcal{M}(X):\exists k_{0}\geqslant 1; \nu\in\displaystyle\bigcap_{l=1}^{q_{j^{\prime},k}}\bigcup_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{Q}_{\tau_{j^{\prime},k}^{l}(i)}\ \text{for all}\ k\geqslant k_{0}\Big\}.$$
    		Clearly, $\widetilde{\mathcal{H}}=\widetilde{\Lambda}\cup\widetilde{Q}_{j^{\prime}}$ and so, $(\widetilde{\mathcal{H}}\cap\widetilde{\mathcal{K}})\setminus\widetilde{Q}_{j^{\prime}}\subseteq\widetilde{\Lambda}\cap\widetilde{\mathcal{K}}$. Since $\widetilde{Q}_{j^{\prime}}$ is closed, we see that every $\nu\in\mathcal{M}(supp\ \mu)\setminus\widetilde{Q}_{j^{\prime}}$ can be arbitrarily well-approximated by elements from $\widetilde{\Lambda}\cap\widetilde{\mathcal{K}}\cap\mathcal{M}(supp\ \mu)$. 
    		
    		More precisely, we have that for every $\nu\in\mathcal{M}(supp\ \mu)\setminus\widetilde{Q}_{j^{\prime}}$ and every $\varepsilon>0$ there are $k_{0}\geqslant 1$ and $\nu^{\prime}\in\mathcal{M}(supp\ \mu)$ satisfying:
    		\begin{align*}&\nu^{\prime}\in\displaystyle\bigcap_{l=1}^{q_{j^{\prime},k}}\bigcup_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{Q}_{\tau_{j^{\prime},k}^{l}(i)} \ \text{for all}\
    			k\geqslant k_{0},\\
    			&\nu^{\prime}(N)=0\ \text{and}\\
    			&\nu^{\prime}(B)\leqslant \nu(B^{\varepsilon})+\varepsilon \  \text{for all Borel sets} \ B\subseteq X.
    		\end{align*}
    		Without loss of generality we may assume that $j^{\prime}=0$. Put $j^{\prime\prime}:=\{0,1\}\setminus\{j^{\prime}\}$; so, $j^{\prime\prime}=1$. Thus, we can rewrite the above condition as follows:
    		
    		$(*)$ For every $\nu\in\mathcal{M}(supp\ \mu)$ such that $\nu(\overline{A})<\eta$ and every $\varepsilon>0$ there are $k_{0}\geqslant 1$ and  $\nu^{\prime}\in\mathcal{M}(supp\ \mu)$ satisfying:
    		\begin{align*}&\nu^{\prime}\in\displaystyle\bigcap_{l=1}^{q_{j^{\prime},k}}\bigcup_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{Q}_{\tau_{j^{\prime},k}^{l}(i)} \ \text{for all}\ k\geqslant k_{0},\\
    			&\nu^{\prime}(N)=0\ \text{and}\\
    			&\nu^{\prime}(B)\leqslant \nu(B^{\varepsilon})+\varepsilon \  \text{for all Borel sets} \ B\subseteq X.
    		\end{align*}
    		On the other hand, since $\displaystyle\widetilde{P}_ {j^{\prime}}\subseteq\bigcup_{l=1}^{q_{j^{\prime},k}}\bigcap_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{j^{\prime},k}^{l}(i)}$ implies $\displaystyle P_{j^{\prime}}\subseteq\bigcup_{l=1}^{q_{j^{\prime},k}}\bigcap_{i=1}^{n_{k}}T^{-i}P_{\sigma_{j^{\prime},k}^{l}(i)}$, it follows from (\ref{notapproximatedincluded}) that $\displaystyle\mu\big(\bigcup_{l=1}^{q_{j^{\prime},k}}\bigcap_{i=1}^{n_{k}}T^{-i}P_{\sigma_{j^{\prime},k}^{l}(i)}\setminus P_{j^{\prime}}\big)\geqslant\alpha$ for every $k\geqslant 1$. Therefore, we may pick some $x\in supp\ \mu$, $x\notin N$ such that $x\in\displaystyle\bigcup_{l=1}^{q_{j^{\prime},k}}\bigcap_{i=1}^{n_{k}}T^{-i}P_{\sigma_{j^{\prime},k}^{l}(i)}\setminus P_{j^{\prime}}$ for infinitely many $k's$. Consequently, $\delta_{x}\in\mathcal{M}(supp\ \mu)$ is such that $\delta_{x}\in\displaystyle\bigcup_{l=1}^{q_{j^{\prime},k}}\bigcap_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{j^{\prime},k}^{l}(i)}\cap \widetilde{P}_{j^{\prime\prime}}$ for infinitely many $k's$. Furthermore, since $x\in P_{1}=A^{c}$ and $x\notin\partial A^{c}$, we have $x\notin\overline{A}$, which implies $\delta_{x}\in\mathcal{M}(supp\ \mu)\setminus\widetilde{Q}_{0}$. Therefore, by $(*)$ above with $\nu=\delta_{x}$ we see that for every $\varepsilon>0$ there are $k_{0}\geqslant 1$ and some $\nu^{\prime}\in\mathcal{M}(supp\ \mu)$ such that:
    		\begin{align}&\nu^{\prime}\in\displaystyle\bigcap_{l=1}^{q_{j^{\prime},k}}\bigcup_{i=1}^{n_{k}}\widetilde{T}^{-i}\widetilde{Q}_{\tau_{j^{\prime},k}^{l}(i)} \ \text{for all}\ k\geqslant k_{0}\label{nuprimein},\\
    			&\nu^{\prime}(N)=0\label{nuprimenull}\ \text{and}\\
    			&\nu^{\prime}(B)\leqslant \delta_{x}(B^{\varepsilon})+\varepsilon \  \text{for all Borel sets} \ B\subseteq X\label{nuprimedeltax}.
    		\end{align}
    		Now, fix some $k^{\prime}\geqslant k_{0}$ such that $\delta_{x}\in\displaystyle\bigcup_{l=1}^{q_{j^{\prime},k^{\prime}}}\bigcap_{i=1}^{n_{k^{\prime}}}\widetilde{T}^{-i}\widetilde{P}_{\sigma_{j^{\prime},k^{\prime}}^{l}(i)}\cap \widetilde{P}_{1}$. So, there exists some $1\leqslant l^{\prime}\leqslant q_{j^{\prime},k^{\prime}}$ such that $\widetilde{T}^{i}\delta_{x}\in\widetilde{P}_{\sigma_{j^{\prime},k^{\prime}}^{l^{\prime}}(i)}$ for every $1\leqslant i\leqslant n_{k^{\prime}}$. That is, we have $\delta_{T^{i}x}(A)>\eta$ if $\sigma_{j^{\prime},k^{\prime}}^{l^{\prime}}(i)=0$ and $\delta_{T^{i}x}(A^{c})\geqslant 1-\eta$ if $\sigma_{j^{\prime},k^{\prime}}^{l^{\prime}}(i)=1$ ($1\leqslant i\leqslant n_{k^{\prime}}$). On the other hand, by \eqref{nuprimein} we see that for every $1\leqslant l\leqslant q_{j^{\prime},k^{\prime}}$ there exists some $1\leqslant i^{\prime}\leqslant n_{k^{\prime}}$ such that $\widetilde{T}^{i^{\prime}}\nu^{\prime}\in\widetilde{Q}_{\tau_{j^{\prime},k^{\prime}}^{l}(i^{\prime})}$. In particular, for $l=l^{\prime}$ there exists some $1\leqslant i^{\prime}\leqslant n_{k^{\prime}}$ such that $\nu^{\prime}(T^{-i^{\prime}}\overline{A})\geqslant \eta$ if $\tau_{j^{\prime},k^{\prime}}^{l^{\prime}}(i^{\prime})=0$ or $\nu^{\prime}(T^{-i^{\prime}}\overline{A^{c}})\geqslant 1-\eta$ if $\tau_{j^{\prime},k^{\prime}}^{l^{\prime}}(i^{\prime})=1$. Now, without loss of generality we may assume that $\sigma_{j^{\prime},k^{\prime}}^{l^{\prime}}(i^{\prime})=0$ (which is the same as $\tau_{j^{\prime},k^{\prime}}^{l^{\prime}}(i^{\prime})=1)$. In this case, we have $T^{i^{\prime}}x\in A$ and $\nu^{\prime}(T^{-i^{\prime}}\overline{A^{c}})\geqslant 1-\eta$. By \eqref{nuprimenull} the condition $\nu^{\prime}(T^{-i^{\prime}}\overline{A^{c}})\geqslant 1-\eta$ is equivalent to $\nu^{\prime}(T^{-i^{\prime}}A^{c})\geqslant 1-\eta$. Moreover, by \eqref{nuprimedeltax} with $B=T^{-i^{\prime}}A^{c}$ we get $\delta_{x}((T^{-i^{\prime}}A^{c})^{\varepsilon})\geqslant \nu^{\prime}(T^{-i^{\prime}}A^{c})-\varepsilon\geqslant 1-\eta-\varepsilon>0$, whenever $\varepsilon>0$ is small enough depending on $0<\eta<1$. Therefore, we see that $x\in T^{-i^{\prime}}A$ and $x\in [(T^{-i^{\prime}}A)^{c}]^{\varepsilon}$ for every $\varepsilon>0$ small enough depending on $0<\eta<1$. Finally, since $i^{\prime}=i^{\prime}(\varepsilon)$ we have to consider two cases:
    		\begin{itemize}
    			\item[(i)]The set $\{i^{\prime}(\varepsilon):\varepsilon>0\}$ is bounded as $\varepsilon\rightarrow 0$.
    			In this case, there are $i^{\prime}\geqslant 1$ and a sequence $\varepsilon_{n}\rightarrow 0$ such that $i^{\prime}=i^{\prime}(\varepsilon_{n})$ for all $n\geqslant 1$. So, we get $x\in T^{-i^{\prime}}A$ and $x\in [(T^{-i^{\prime}}A)^{c}]^{\varepsilon_{n}}$ for all $n\geqslant 1$. Therefore, by letting $n\rightarrow\infty$ we conclude that $x\in T^{-i^{\prime}}A\ \cap\overline{(T^{-i^{\prime}}A)^{c}}$, which contradicts the choice $x\notin N$.
    			\item[(ii)]The set $\{i^{\prime}(\varepsilon):\varepsilon>0\}$ is unbounded as $\varepsilon\rightarrow 0$. In this case, by letting $\varepsilon\rightarrow 0$ we obtain $\displaystyle x\in\overline{\bigcup_{r\in\mathbb{Z}}T^{r}A}\ \cap \overline{\bigcup_{r\in\mathbb{Z}}\partial T^{r}A^{c}}$, which contradicts the choice $x\notin N$ again.
    		\end{itemize}
    		This concludes the proof of the theorem.
    	\end{proof}
    \end{theorem}
    
    \begin{corollary}\label{Corollary1} Let $(X,\mathcal{X},\mu,T)$ be a MDS and let $\widetilde{\mu}\in Q(\mu)$ having full-support.
    	The following property holds:
    	
    	Given $0<\alpha<1$ there exists $\beta>0$ such that, if $\displaystyle\widetilde{\mathcal{P}}\subseteq_{\beta}^{\widetilde{\mu}}\bigvee_{i=1}^{\infty}\widetilde{T}^{-i}\widetilde{\mathcal{P}}$, then $\displaystyle\mathcal{P}\subseteq_{\alpha}^{\mu}\bigvee_{i=1}^{\infty}T^{-i}\mathcal{P}$.
    	\begin{proof} Let $0<\alpha<1$ be arbitrary and take $\beta>0$ and $n_{0}\geqslant 1$ as in Theorem~\ref{approximatedbypast}. Suppose $\displaystyle\widetilde{\mathcal{P}}\subseteq_{\beta}^{\widetilde{\mu}}\bigvee_{i=1}^{\infty}\widetilde{T}^{-i}\widetilde{\mathcal{P}}$. There exists $n_{1}\geqslant 1$ such that $\displaystyle\widetilde{\mathcal{P}}\subseteq_{\beta}^{\widetilde{\mu}}\bigvee_{i=1}^{n}\widetilde{T}^{-i}\widetilde{\mathcal{P}}$, whenever $n\geqslant n_{1}$. Now, fix any $n\geqslant\max\{n_{0},n_{1}\}$. Since $\displaystyle\widetilde{\mathcal{P}}\subseteq_{\beta}^{\widetilde{\mu}}\bigvee_{i=1}^{n}\widetilde{T}^{-i}\widetilde{\mathcal{P}}$ whenever $n\geqslant n_{1}$, by Theorem~\ref{approximatedbypast} we conclude that $\displaystyle\mathcal{P}\subseteq_{\alpha}^{\mu}\bigvee_{i=1}^{n}T^{-i}\mathcal{P}$. Since $\displaystyle\bigvee_{i=1}^{n}T^{-i}\mathcal{P}\subseteq\bigvee_{i=1}^{\infty}T^{-i}\mathcal{P}$ we get $\displaystyle\mathcal{P}\subseteq\bigvee_{i=1}^{\infty}T^{-i}\mathcal{P}$, as desired.
    	\end{proof}
    \end{corollary}
    
    \begin{theorem}\label{zeroentropy}Let $(X,\mathcal{X},\mu,T)$ be a MDS and let $\widetilde{\mu}\in Q(\mu)$ having full-support. If $h_{\widetilde{\mu}}(\widetilde{T},\widetilde{\mathcal{P}})=0$, then $h_{\mu}(T,\mathcal{P})=0$.
    	
    	\begin{proof} Suppose $h_{\widetilde{\mu}}(\widetilde{T},\widetilde{P})=0$ and let $0<\alpha<1$ be arbitrary. Pick $\beta>0$ as in Corollary~\ref{Corollary1}. Since $h_{\widetilde{\mu}}(\widetilde{T},\widetilde{P})=0$, we have $\displaystyle\widetilde{\mathcal{P}}\subseteq_{\beta}^{\widetilde{\mu}}\bigvee_{i=1}^{\infty}\widetilde{T}^{-i}\widetilde{\mathcal{P}}$ and so, by Corollary~\ref{Corollary1} we get $\displaystyle\mathcal{P}\subseteq_{\alpha}^{\mu}\bigvee_{i=1}^{\infty}T^{-i}\mathcal{P}$. Since $0<\alpha<1$ is arbitrary, we obtain $\displaystyle\mathcal{P}\subseteq_{0}^{\mu}\bigvee_{i=1}^{\infty}T^{-i}\mathcal{P}$, which is equivalent to $h_{\mu}(T,\mathcal{P})=0$.
    	\end{proof}
    \end{theorem}
    
    \smallskip

    Finally, we are ready to prove our main result.
    \begin{theorem}{\label{Mainresult}} Let $(X,\mathcal{X},\mu,T)$ be a MDS with $h_{\mu}(T)<\infty$. If $h_{\mu}(T)>0$, then $h_{\widetilde{\mu}}(\widetilde{T})>0$ for every $\widetilde{\mu}\in Q(\mu)$ having full-support.
    	
    	\begin{proof} Suppose $h_{\mu}(T)>0$ and let $\widetilde{\mu}\in Q(\mu)$ having full-support. By the Downarowicz-Huczek theorem \cite{Downarowicz_Huczek}, there exist a zero-dimensional extension $\pi:Y\rightarrow X$ and a unique Borel probability measure $\nu$ on $Y$ which is invariant under the shift $S$ such that $\mu=\nu\circ \pi^{-1}$ and $h_{\nu}(S)=h_{\mu}(T)>0$. So, there exists a clopen set $B\subset Y$ with $h_{\nu}(S,\mathcal{Q})>0$, where $\mathcal{Q}=\{B,B^{c}\}$. Given any $0<\eta<1$, if we put $\displaystyle\widetilde{\mathcal{Q}}=\{\widetilde{B},\widetilde{B}^{c}\}$, where $\widetilde{B}=\{\theta:\theta(B)>\eta\}$, by Theorem~\ref{zeroentropy} it follows that $h_{\widetilde{\nu}}(\widetilde{S},\widetilde{\mathcal{Q}})>0$, where $\widetilde{\nu}:=\pi^{-1}\widetilde{\mu}$ is a quasifactor of $\nu$ having full-support . Therefore, we obtain $h_{\widetilde{\mu}}(\widetilde{T},\pi\widetilde{Q})>0$, which implies $h_{\widetilde{\mu}}(\widetilde{T})>0$, as desired.
    	\end{proof}
    \end{theorem}
    
    \section*{Acknowledgments}
    	The author would like to dedicate this work to his daughter, Laura C. Vermersch, in gratitude for the joyful moments during its writing.

    %Equations in \LaTeX{} can either be inline or on-a-line by itself. For
    %inline equations use the \verb+$...$+ commands. Eg: The equation
    %$H\psi = E \psi$ is written via the command $H \psi = E \psi$.
    
    %For on-a-line by itself equations (with auto generated equation numbers)
    %one can use the equation or eqnarray environments \textit{D}.
    %\begin{equation}
    %\mathcal{L} = i {\psi} \gamma^\mu D_\mu \psi
    %  - \frac{1}{4} F_{\mu\nu}^a F^{a\mu\nu} - m {\psi} \psi
    %\label{eq1}
    %\end{equation}
    %where,
    %\begin{align}
    %D_\mu &=  \partial_\mu - ig \frac{\lambda^a}{2} A^a_\mu
    %\nonumber \\
    %F^a_{\mu\nu} &= \partial_\mu A^a_\nu - \partial_\nu A^a_\mu
    %  + g f^{abc} A^b_\mu A^a_\nu
    %\label{eq2}
    %\end{align}
    %Notice the use of \verb+\nonumber+ in the align environment at the end
    %of each line, except the last, so as not to produce equation numbers on
    %lines where no equation numbers are required. The \verb+\label{}+ %command
    %should only be used at the last line of an align environment where
    %\verb+\nonumber+ is not used.
    %\begin{equation}
    %Y_\infty = \left( \frac{m}{\textrm{GeV}} \right)^{-3}
    %  \left[ 1 + \frac{3 \ln(m/\textrm{GeV})}{15}
    % + \frac{\ln(c_2/5)}{15} \right]
    %\end{equation}
    %The class file also supports the use of \verb+\mathbb{}+, \verb+\mathscr{}+ %and
    %\verb+\mathcal{}+ commands. As such \verb+\mathbb{R}+, %\verb+\mathscr{R}+
    %and \verb+\mathcal{R}+ produces $\mathbb{R}$, $\mathscr{R}$ and %$\mathcal{R}$
    %respectively.

   %	\section*{Acknowledgments}
    
    	%	The author would like to dedicate this work to his daughter, Laura C. %Vermersch, in gratitude for the joyful moments during its writing.

   % \end{Backmatter}
    
    \end{document}